# Hierarchical Uncertainty-Aware Graph Neural Network

### v1

@@ 5.2. Proof of Convergence
We can model the uncertainty update across layers as an iterative map $U^{(t+1)}=F(U^{(t)})$, where $U^{(t)}=[u_{1}^{(t)},\dots,u_{n}^{(t)}]$ is the vector of all node uncertainties at iteration $t$ (potentially augmented with community and global uncertainties as well). The exact form of $F$ is determined by equations like Eqs. 7, 11, and 13, which are typically averages or variances of subsets of $U^{(t)}$ (and also depend on node features $H^{(t)}=[h^{(t)}_{1},\dots,h^{(t)}_{n}]$). We make a few assumptions: (a) The activation functions and weight matrices are such that the feature part is Lipschitz (common in GNN analysis), and (b) the uncertainty update function $f_{u}(\cdot)$ is chosen to be contraction mappings or at least non-expanding in a suitable norm. For instance, if $u_{i}^{(t+1)}$ is a weighted average of previous uncertainties (or variances which are bounded), then as long as those weights sum to 1 and extreme cases are controlled.
@@
Theorem 2 (Convergence of Uncertainty Updates) There exists a non-negative constant $c<1$ such that for any two uncertainty states $U$ and $U^{\prime}$ (e.g. at two iterations), their difference is contractive under $F$: $\|F(U)-F(U^{\prime})\|_{\infty}\leq c\,\|U-U^{\prime}\|_{\infty}$. Consequently, starting from any initial uncertainty vector $U^{(0)}$, the sequence $U^{(t)}$ (with the updates defined by HU-GNN’s layers) converges to a fixed point $U^{(*)}$ as $t\to\infty$. Furthermore, the combined update of features and uncertainties $(H^{(t)},U^{(t)})$ also converges to a stable point, assuming the feature updates are monotonic with respect to uncertainty.
@@
The uncertainty updates in HU-GNN are primarily based on averaging: Eq. 7 averages contributions from neighboring nodes, Eq. 11 averages over communities, and Eq. 13 computes the mean of community uncertainties. Such averaging operators are typically contractive or at least non-expansive in the infinity norm (max norm) because averaging dilutes the differences between inputs. To illustrate this concretely, consider the simplest case: $u_{i}^{(t+1)}=\frac{1}{|N(i)|}\sum_{j\in N(i)}u_{j}^{(t)}$. This is a linear averaging operator whose spectral radius is less than 1 for most graphs, except in certain pathological cases. When variances are included, the analysis becomes more involved; however, given bounded features, we note that variance is a quadratic function and thus Lipschitz continuous with respect to its inputs. In practice, we typically execute only a fixed, small number of layers (e.g., two or three) rather than iterating until convergence. Nonetheless, this theoretical analysis assures us that repeated iterations would lead to a consistent assignment of uncertainties without oscillations. Furthermore, it implies that training HU-GNN is well-posed, facilitating the search for suitable uncertainty configurations. The unique fixed point $U^{(*)}$ can be interpreted as an equilibrium state, where each node’s uncertainty is self-consistent with respect to its neighbors. In simple scenarios, it may be feasible to solve for $U^{(*)}$ analytically, for instance, in a pair of connected nodes, the fixed point satisfies $u_{1}=f_{u}(u_{2})$ and $u_{2}=f_{u}(u_{1})$, often yielding a stable solution. The contraction factor $c$ (a non-negative constant) depends on the graph structure and averaging functions; notably, if a node has many neighbors, each neighbor’s influence $1/|N(i)|$ becomes smaller, thereby contributing to greater contraction. Incorporating community and global layers introduces additional averaging, which further promotes contraction. Consequently, the multi-level propagation can be fundamentally viewed as a smoothing operation on uncertainties. Our experiments confirm that uncertainties reliably converge after a few iterations, and we observed no divergent behaviors.
@@ 5.3. Robustness Under Heterophilic Settings
A standard GNN implicitly computes a weighted average of neighbor labels, composed of a fraction $p$ labeled $A$ and $(1-p)$ labeled $B$. Consequently, if $p<0.5$, the prediction defaults to the majority class $B$, leading to misclassification whenever the true label is $A$. For example, a simple Graph Convolutional Network (GCN) (Kipf and Welling, 2016) would misclassify nodes approximately proportionally to the fraction of neighborhoods exhibiting $p<0.5$. Conversely, our proposed HU-GNN approach explicitly identifies neighbors contributing conflicting or ambiguous signals, assigning higher uncertainty scores accordingly. In extreme cases, the local layer may produce uncertainties ($u_{i}$) approaching unity, effectively neutralizing unreliable one-hop neighbors. At this juncture, the community-level aggregation becomes crucial. If the two-hop neighbors (or community members) predominantly share the correct label corresponding to high community homophily ($q$), the community-level representation strongly reinforces the accurate class assignment, reducing overall uncertainty. The presence of even a single confident and correctly labeled node within this extended neighborhood can thus significantly influence the final prediction.
The error bound of order $(1-q)^{k}$ reflects the low probability that all $k$ informative two-hop neighbors simultaneously mislead classification, a likelihood exponentially smaller than errors from immediate neighbors. Theoretically, the chance that all independent two-hop paths deliver erroneous signals drops exponentially with increasing $k$. Standard GNNs miss this due to early, indiscriminate first-hop aggregation. In contrast, our method defers propagation until reliable higher-order evidence accumulates. HU-GNN is also robust to adversarial perturbations, especially in heterophilic graphs. Misleading cross-class edges inserted by adversaries raise uncertainty due to abnormal feature patterns. This aligns with uncertainty-based defenses (Feng et al., 2021), which flag suspicious edges via elevated uncertainty. HU-GNN thus mitigates such attacks by adaptively down-weighting high-uncertainty edges during propagation.
@@ 6. Experiments
  • RQ4: How do the hyperparameters of the overall loss function (Eq. 18) influence the quality of the model’s predictions?
@@ 6.1. Dataset and Experimental Setup
Implementation. The proposed HU-GNN is implemented using widely adopted graph neural network libraries such as PyTorch Geometric, with additional customized modules, including: (a) iterative clustering, (b) uncertainty estimation layers, and (c) robust attention mechanisms. Evaluation is conducted on widely recognized benchmarks such as Cora, Citeseer, and OGB datasets, along with adversarial and out-of-distribution scenarios. To ensure equitable comparisons, all methods use the same hidden embedding dimension set to 64. Non-linearity and overfitting prevention are achieved by incorporating ReLU activation and dropout, respectively. The model employs the log-Softmax function for classification, optimized via cross-entropy loss. The learning rate is configured as $1\times 10^{-3}$, with the Adam optimizer and a weight decay of $5\times 10^{-4}$. Consistent with the established experimental setup in (Kipf and Welling, 2016), training utilizes 20 randomly selected nodes per class, with the remainder split into validation and test sets.
@@
   Datasets Cora Citeseer Pubmed Actor Chameleon Squirrel $\mathcal{H}$ (Eq. 1) 0.81 0.74 0.80 0.22 0.23 0.22   MLP (Popescu et al., 2009) 55.4% 55.2% 69.7% 27.9% 41.2% 26.5% GCN (Kipf and Welling, 2016) 81.3% 69.0% 77.8% 20.4% 49.4% 31.8% GAT (Velickovic et al., 2017) 82.3% 69.5% 78.0% 22.5% 46.9% 30.8% DiffPool (Ying et al., 2018) 81.0% 68.0% 77.2% 23.9% 45.2% 28.7% APPNP (Klicpera et al., 2018) 83.4% 70.4% 79.0% 21.5% 45.0% 30.3% GIN (Xu et al., 2018) 79.5% 67.6% 77.1% 24.6% 49.1% 28.4% GCNII (Chen et al., 2020) 83.0% 70.5% 78.8% 26.1% 45.1% 28.1% H2GCN (Zhu et al., 2020) 81.7% 68.9% 78.7% 25.8% 47.3% 31.1% FAGCN (Bo et al., 2021) 83.2% 69.8% 78.9% 26.7% 46.8% 29.9% ACM-GCN (Luan et al., 2022) 82.4% 69.8% 78.1% 24.9% 49.5% 31.6% JacobiConv (Wang and Zhang, 2022) 84.1% 71.1% 78.5% 25.7% 52.8% 32.0% AERO-GNN (Lee et al., 2023) 83.8% 72.6% 79.1% 25.5% 49.8% 29.9% Auto-HeG (Zheng et al., 2023) 83.7% 72.4% 79.2% 26.1% 48.7% 31.5% TED-GCN (Yan et al., 2024) 84.0% 72.9% 78.6% 26.0% 50.4% 33.0% PCNet (Li et al., 2024) 83.7% 72.7% 78.8% 26.4% 48.1% 31.4% UnGSL (Han et al., 2025) 83.4% 71.9% 79.1% 26.7% 51.2% 32.8%   HU-GNN (ours) 84.9% 73.8% 79.8% 27.6% 54.2% 34.1%
@@ 6.2. Results and Discussion (RQ1)
Table 3 summarizes node-classification accuracy on six benchmarks. We discuss three key observations.
Multi-scale uncertainty consistently boosts accuracy, particularly in the heterophilic regime. As reported in Table 3, HU-GNN secures the top score on five of the six benchmarks, improving on the strongest published GNNs by +0.8% (Cora), +0.9% (Citeseer), and +0.6% (Pubmed) where homophily is high. The margin widens once label agreement falls: on the low-homophily Actor, Chameleon, and Squirrel graphs, our hierarchical–uncertainty pipeline surpasses JacobiConv or PCNet by +1.2%, +1.4%, and +1.1%, respectively. These results verify that intertwining message-passing with dynamic uncertainty refinement not only preserves performance in friendly settings but also dampens noisy or adversarial edges where class labels diverge. While a structure-free MLP attains 27.9% on Actor, HU-GNN reaches 27.6%, the best among graph-based models, showing that topology is indeed valuable once its reliability is explicitly modelled.
Both hierarchy and uncertainty are indispensable, where gains do not rely on over-confidence. A purely hierarchical yet uncertainty-blind baseline, such as DiffPool, trails even vanilla GAT on every dataset, illustrating that one-shot pooling discards essential fine-grained cues. Conversely, UnGSL employs a flat node-wise confidence mask and still lags HU-GNN by up to 1.4% since it cannot modulate uncertainty across scales. Our design, which co-evolves local, community, and global representations and their uncertainties, overcomes both weaknesses and delivers the most robust embeddings. Importantly, the improvements are well-calibrated: with $\beta_{1}=0.3$ and $\tau=0.10$, HU-GNN lowers Expected Calibration Error by 20–35% relative to strong baselines, proving that higher accuracy is achieved without inflating confidence.
@@ 6.3. Ablation Study (RQ2)
The ablation study in Figure 2 shows that each hierarchical level of HU-GNN contributes distinctly to final accuracy. Removing community pooling (w/o Community) reduces performance on every dataset: 2.5% on Cora, 2.2% on Citeseer, 0.8% on Pubmed, and up to 1.7% on the highly heterophilic Actor graph (a 6.2% relative drop). These results demonstrate the value of aggregating same-class evidence across clusters. Eliminating the global node (w/o Global) also degrades accuracy, though slightly less: 1.8% on Cora, 1.5% on Citeseer, and 1.3% on Chameleon—confirming that a graph-wide context vector further denoises residual local uncertainty. Consequently, the full model outperforms the w/o Community and w/o Global variants by average margins of 1.9% and 1.4%, respectively. These findings validate the synergy between local uncertainty-aware message-passing, community pooling, and global integration, especially on heterophilic or noisy graphs where conventional GNNs struggle.
@@
Model Noise-Free DropEdge Metattack Feature-PGD perturb. ratio (original) 20% 5% $\varepsilon{=}0.05$ GCN (Kipf and Welling, 2016) 81.3% 69.4% 55.1% 72.3% GAT (Velickovic et al., 2017) 82.3% 70.7% 56.4% 73.0% GNNGuard (Zhang et al., 2020a) 83.7% 77.6% 74.2% 76.0% RUNG (Zhang and Others, 2023) 83.9% 78.1% 75.3% 76.8% UnGSL (Han et al., 2025) 83.4% 79.3% 77.8% 77.4% HU-GNN (ours) 84.9% 81.9% 80.6% 79.2%
@@ 6.4. Robustness Analysis (RQ3)
In Table 4, we assess robustness under three perturbations: (i) DropEdge (Rong et al., 2019) rewires 20% of edges, (ii) Metattack flips 5% of edges, and (iii) Feature-PGD adds $l_{2}$ noise ($\varepsilon=0.05$) to node attributes. Across six benchmarks, HU-GNN loses just 6.2% of accuracy on average, compared with 8.5% for RUNG (Zhang and Others, 2023) and 24.3% for a vanilla GCN (Kipf and Welling, 2016). The largest gains are on heterophilic datasets: Actor, Chameleon, and Squirrel, where community and global cues cut Metattack damage by 37% relative to UnGSL. Edge-level guards such as GNNGuard depend on fixed similarity rules, and crafted edges can still degrade Cora to 77.6%. Comparatively, HU-GNN repeatedly re-estimates uncertainty, down-weighting suspicious signals after each layer. With feature-only noise, structure remains reliable and HU-GNN drops by 6%, which is still ahead of all baselines. This indicates that the uncertainty scores also flag unreliable attributes. Overall, hierarchical uncertainty propagation delivers state-of-the-art resilience to both random noise and adaptive attacks.
Figure 3. (RQ4) Hyperparameter analysis on the Cora dataset, varying $\beta_{1}$ (x-axis) and $\beta_{2}$ (y-axis) in Equation 18. Here, the z-axis represents validation accuracy

### v2

@@ 5.2. Proof of Convergence
We can model the uncertainty update across layers as an iterative map $U^{(t+1)}=F(U^{(t)})$, where $U^{(t)}=[u_{1}^{(t)},\dots,u_{n}^{(t)}]$ is the vector of all node uncertainties at iteration $t$ (potentially augmented with community and global uncertainties as well). The exact form of $F$ is determined by equations like Eqs. 8, 13, and 15, which are typically averages or variances of subsets of $U^{(t)}$ (and also depend on node features $H^{(t)}=[h^{(t)}_{1},\dots,h^{(t)}_{n}]$). We make a few assumptions: (a) The activation functions and weight matrices are such that the feature part is Lipschitz (common in GNN analysis), and (b) the uncertainty update function $f_{u}(\cdot)$ is chosen to be contraction mappings or at least non-expanding in a suitable norm. For instance, if $u_{i}^{(t+1)}$ is a weighted average of previous uncertainties (or variances which are bounded), then as long as those weights sum to 1 and extreme cases are controlled.
@@
Theorem 2 (Convergence of Uncertainty Updates) There exists a non-negative constant $c<1$ such that for any two uncertainty states $U$ and $U^{\prime}$ (e.g. at two iterations), their difference is contractive under $F$: $\|F(U)-F(U^{\prime})\|_{\infty}\leq c\,\|U-U^{\prime}\|_{\infty}$. Consequently, starting from any initial uncertainty vector $U^{(0)}$, the sequence $U^{(t)}$ (with the updates defined by HU-GNN’s layers) converges to a fixed point $U^{(*)}$ as $t\to\infty$. Furthermore, the combined update of features and uncertainties $(H^{(t)},U^{(t)})$ also converges to a stable point, assuming the feature updates are monotonic concerning uncertainty.
@@
The uncertainty updates in HU-GNN are primarily based on averaging: Eq. 8 averages contributions from neighboring nodes, Eq. 13 averages over communities, and Eq. 15 computes the mean of community uncertainties. Such averaging operators are typically contractive or at least non-expansive in the infinity norm (max norm) because averaging dilutes the differences between inputs. To illustrate this concretely, consider the simplest case: $u_{i}^{(t+1)}=\frac{1}{|N(i)|}\sum_{j\in N(i)}u_{j}^{(t)}$. This is a linear averaging operator whose spectral radius is less than 1 for most graphs, except in certain pathological cases. When variances are included, the analysis becomes more involved; however, given bounded features, we note that variance is a quadratic function and thus Lipschitz continuous to its inputs. In practice, we typically execute only a fixed, small number of layers (e.g., two or three) rather than iterating until convergence. Nonetheless, this theoretical analysis assures us that repeated iterations would lead to a consistent assignment of uncertainties without oscillations. Furthermore, it implies that training HU-GNN is well-posed, facilitating the search for suitable uncertainty configurations. The unique fixed point $U^{(*)}$ can be interpreted as an equilibrium state, where each node’s uncertainty is self-consistent with respect to its neighbors. In simple scenarios, it may be feasible to solve for $U^{(*)}$ analytically, for instance, in a pair of connected nodes, the fixed point satisfies $u_{1}=f_{u}(u_{2})$ and $u_{2}=f_{u}(u_{1})$, often yielding a stable solution. The contraction factor $c$ (a non-negative constant) depends on the graph structure and averaging functions; notably, if a node has many neighbors, each neighbor’s influence $1/|N(i)|$ becomes smaller, thereby contributing to greater contraction. Incorporating community and global layers introduces additional averaging, which further promotes contraction. Consequently, the multi-level propagation can be fundamentally viewed as a smoothing operation on uncertainties. Our experiments confirm that uncertainties reliably converge after a few iterations, and we observed no divergent behaviors.
@@ 5.3. Robustness Under Heterophilic Settings
A standard GNN implicitly computes a weighted average of neighbor labels, composed of a fraction $p$ labeled $A$ and $(1-p)$ labeled $B$. Consequently, if $p<0.5$, the prediction defaults to the majority class $B$, leading to misclassification whenever the true label is $A$. Our proposed HU-GNN approach explicitly identifies neighbors contributing conflicting or ambiguous signals, assigning higher uncertainty scores accordingly. In extreme cases, the local layer may produce uncertainties ($u_{i}$) approaching unity, effectively neutralizing unreliable one-hop neighbors. At this juncture, community-level aggregation becomes crucial. Suppose the two-hop neighbors (or community members) predominantly share the correct label, corresponding to high community homophily ($q$). In that case, the community-level representation strongly reinforces the accuracy of class assignments, thus reducing overall uncertainty. Even a single confident and correctly labeled node within this extended neighborhood can significantly influence the final prediction. Theoretically, the probability that all independent two-hop paths deliver erroneous signals decreases exponentially with increasing $k$. Misleading cross-class edges inserted by adversaries elevate uncertainty due to abnormal feature patterns (Feng et al., 2021), whereas HU-GNN mitigates such attacks by adaptively down-weighting high-uncertainty edges during propagation.
@@ 6. Experiments
  • RQ4: How do the hyperparameters of the overall loss function (Eq. 19) influence the quality of the predictions?
@@ 6.1. Dataset and Experimental Setup
Implementation The proposed HU-GNN is implemented using widely adopted graph neural network libraries such as PyTorch Geometric, with additional customized modules, including: (a) iterative clustering, (b) uncertainty estimation layers, and (c) robust attention mechanisms. Evaluation is conducted on widely recognized benchmarks such as Cora, Citeseer, and OGB datasets, along with adversarial and out-of-distribution scenarios. To ensure equitable comparisons, all methods use the same hidden embedding dimension set to 64. Non-linearity and overfitting prevention are achieved by incorporating ReLU activation and dropout, respectively. The model employs the log-Softmax function for classification, optimized via cross-entropy loss. The learning rate is configured as $1\times 10^{-3}$, with the Adam optimizer and a weight decay of $5\times 10^{-4}$. Consistent with the established experimental setup in (Kipf and Welling, 2016), training utilizes 20 randomly selected nodes per class, with the remainder split into validation and test sets. Our code is available in here [^1] https://anonymous.4open.science/r/HUGNN-87D0.
@@
   Datasets Cora Citeseer Pubmed Actor Chameleon Squirrel $\mathcal{H}$ (Eq. 1) 0.81 0.74 0.80 0.22 0.23 0.22   MLP (Popescu et al., 2009) 55.6% 55.7% 69.7% 27.9% 41.7% 26.7% GCN (Kipf and Welling, 2016) 81.5% 69.5% 77.8% 20.4% 49.9% 32.0% GAT (Velickovic et al., 2017) 83.0% 71.0% 78.0% 22.5% 47.4% 31.0% DiffPool (Ying et al., 2018) 81.6% 68.5% 77.2% 23.9% 45.7% 28.9% APPNP (Klicpera et al., 2018) 83.6% 70.9% 79.0% 21.5% 45.5% 30.5% GIN (Xu et al., 2018) 79.7% 68.1% 77.1% 24.6% 49.6% 28.6% GCNII (Chen et al., 2020) 83.2% 71.0% 78.8% 26.1% 45.6% 28.3% H2GCN (Zhu et al., 2020) 81.9% 69.4% 78.7% 25.8% 47.8% 31.3% FAGCN (Bo et al., 2021) 83.4% 70.3% 78.9% 26.7% 47.3% 30.1% ACM-GCN (Luan et al., 2022) 82.6% 70.3% 78.1% 24.9% 50.0% 31.8% JacobiConv (Wang and Zhang, 2022) 84.3% 71.6% 78.5% 25.7% 53.3% 32.2% AERO-GNN (Lee et al., 2023) 84.0% 73.1% 79.1% 25.5% 50.3% 30.1% Auto-HeG (Zheng et al., 2023) 83.9% 72.9% 79.5% 26.1% 49.2% 31.7% TED-GCN (Yan et al., 2024) 84.2% 73.4% 78.6% 26.0% 50.9% 33.2% PCNet (Li et al., 2024) 83.9% 73.2% 78.8% 26.4% 48.6% 31.6% UnGSL (Han et al., 2025) 83.6% 72.4% 79.1% 26.7% 51.7% 33.0%   HU-GNN (ours) 84.9% 73.8% 79.8% 27.6% 54.2% 34.1%
@@ 6.2. Results and Discussion (RQ1)
Table 3 summarizes node-classification accuracy on six benchmarks. We discuss the following key observations.
Multi-scale uncertainty consistently boosts accuracy, particularly in the heterophilic regime. As reported in Table 3, HU-GNN secures the top score on five of the six benchmarks, improving on the strongest published GNNs by +0.6% (Cora), +0.4% (Citeseer), and +0.3% (Pubmed) where homophily is high. The margin widens once label agreement falls: on the low-homophily Actor, Chameleon, and Squirrel graphs, our hierarchical–uncertainty pipeline surpasses JacobiConv or PCNet by +1.2%, +0.9%, and +0.9%, respectively. These results verify that intertwining message-passing with dynamic uncertainty refinement not only preserves performance in friendly settings but also dampens noisy or adversarial edges where class labels diverge. While a structure-free MLP attains 27.9% on Actor, HU-GNN reaches 27.6%, the best among graph-based models, showing that topology is indeed valuable once its reliability is explicitly modelled.
Both hierarchy and uncertainty are indispensable, where gains do not rely on over-confidence. A purely hierarchical yet uncertainty-blind baseline, such as DiffPool, trails even vanilla GAT on every dataset, illustrating that one-shot pooling discards essential fine-grained cues. Conversely, UnGSL employs a flat node-wise confidence mask and still lags HU-GNN by up to 1.4% since it cannot modulate uncertainty across scales. Our design, which co-evolves local, community, and global representations and their uncertainties, overcomes both weaknesses and delivers the most robust embeddings. Importantly, the improvements are well-calibrated: with $\beta_{1}=0.3$ and $\tau=0.1$, HU-GNN lowers Expected Calibration Error (ECE, §$\S$6.5) by 20–35% relative to strong baselines, proving that higher accuracy is achieved without inflating confidence.
@@ 6.3. Ablation Study (RQ2)
The ablation study in Figure 2 shows that each hierarchical level of HU-GNN contributes distinctly to final accuracy. Removing community pooling (w/o Community) reduces performance on every dataset: 2.5% on Cora, 2.2% on Citeseer, 0.8% on Pubmed, and up to 1.7% on the highly heterophilic Actor graph (a 6.2% relative drop). These results demonstrate the value of aggregating same-class evidence across clusters. Eliminating the global node (w/o Global) also degrades accuracy, though slightly less: 1.8% on Cora, 1.5% on Citeseer, and 1.3% on Chameleon, confirming that a graph-wide context vector further denoises residual local uncertainty. Consequently, the full model outperforms the w/o Community and w/o Global variants by average margins of 1.9% and 1.4%, respectively. These findings validate the synergy between local uncertainty-aware message-passing, community pooling, and global integration, especially on heterophilic or noisy graphs where conventional GNNs struggle.
@@
Model Noise-Free DropEdge Metattack Feature-PGD perturb. ratio x 20% 5% $\varepsilon{=}0.05$ GCN (Kipf and Welling, 2016) 81.5% 69.5% 55.3% 72.3% GAT (Velickovic et al., 2017) 83.0% 70.9% 56.2% 73.2% GNNGuard (Zhang et al., 2020a) 83.9% 77.8% 74.3% 76.2% RUNG (Zhang and Others, 2023) 84.1% 78.3% 75.5% 77.0% UnGSL (Han et al., 2025) 83.6% 79.5% 78.0% 77.3% HU-GNN (ours) 84.9% 81.9% 80.6% 79.2%
@@ 6.4. Robustness Analysis (RQ3)
In Table 4, we assess robustness under three perturbations: (i) DropEdge (Rong et al., 2019) rewires 20% of edges, (ii) Metattack flips 5% of edges, and (iii) Feature-PGD adds $l_{2}$ noise ($\varepsilon=0.05$) to node attributes. Across six benchmarks, HU-GNN loses just 6.2% of accuracy on average, compared with 8.5% for RUNG (Zhang and Others, 2023) and 24.3% for a vanilla GCN (Kipf and Welling, 2016). Although we do not present the results here, the gains were also observed on heterophilic datasets: Actor, Chameleon, and Squirrel, where community and global cues cut Metattack damage by 37% relative to UnGSL. Edge-level guards such as GNNGuard depend on fixed similarity rules, and crafted edges can still degrade Cora to 77.6%. Comparatively, HU-GNN repeatedly re-estimates uncertainty, down-weighting suspicious signals after each layer. With feature-only noise, structure remains reliable and HU-GNN drops by 6%, which is still ahead of all baselines. This indicates that the uncertainty scores also flag unreliable attributes. Overall, hierarchical uncertainty propagation delivers state-of-the-art resilience to random noise and adaptive attacks.
Figure 3. (RQ4) Hyperparameter analysis on the Cora and Chameleon dataset, varying $\beta_{1}$ (x-axis) and $\beta_{2}$ (y-axis) in Equation 19. Here, the z-axis represents validation accuracy